OpenStudy (nannie):

Please show me how to work this problem! x²+50=10x, to solve, simplify and the answer should be in the form of a +bί. Thanks!

OpenStudy (anonymous):

x = 5 + 5i or x = 5 - 5i

OpenStudy (nannie):

I know the answer, Just want to know how to work the problem

OpenStudy (anonymous):

(x -5 ) ^2 = -25 x -5 = 5i or x-5 = -5i x = 5 + 5i or x = 5 - 5i

OpenStudy (nannie):

How did you get the (x-5)²=-25

OpenStudy (anonymous):

it's called completing the square, a little trick you can read more abotu that here: http://en.wikipedia.org/wiki/Completing_the_square basically what you do is rewrite it to the =0 x^2 -10x +50 then halve the then and put it in brackets with the x and substract 5^2 from the total to compensate (x - 5)^2 +50 - 25

OpenStudy (anonymous):

halve the TEN not the THEN

OpenStudy (nannie):

OK, I remember now! Thank you So much!

OpenStudy (anonymous):

no worries :)